OpenStudy (anonymous):

What is the inverse of g(x)=x-1/4

OpenStudy (anonymous):

A.g^-1(x) = 4x - 4 B.g^-1(x) = 4x - 1 C.g^-1(x) = 4x + 1 D.g^-1(x) = 4x + 4

OpenStudy (anonymous):

\[g(x)=\frac{ x-1 }{ 4 }\]

OpenStudy (shamil98):

y = x - 1 / 4 switch places with x and y x = y - 1 /4 solve for y.

OpenStudy (shamil98):

to get the inverse.

OpenStudy (anonymous):

C.